OpenStudy (anonymous):

1. Use the sequence 3, -6, 12, -24, … to answer the questions below a. Explain, using complete sentences whether this sequence is arithmetic or geometric and why. --- I guessed its geometric but if thats the answer why is it?

OpenStudy (anonymous):

the answer is a

OpenStudy (texaschic101):

in an arithmetic sequence, the next number is found by adding. In a geometric sequence, the next number is found by multiplying. so lets look at your sequence : 3, -6, 12, -24 what has to be done to 3 to get -6....3 x -2 = -6 will that same process work for the next number ? -6 x -2 = -12....and the next...12 x -2 = -24 so the next number was found by multiplying making it a geometric sequence. Now if you would have tried addition.. 3 + ? = -6 3 + (-9) = -6 and if we take that -9 and add it to the next number, do we get the number after that ? -6 + (-9) = -6 - 9 = - 15...not the next number....so it is not arithmetic. Understand ?

OpenStudy (texaschic101):

oops...there is a typo in there....-6 x -2 = 12